# Stability of the expanding region of Kerr de Sitter spacetimes

### v1

@@ 2.2.2 First and second variation equations for differences
$\begin{split}\Phi\operatorname{Ric}_{i}{}^{j}-\widetilde{\Phi}\widetilde{\operatorname{Ric}}_{i}{}^{j}=&\phantom{+}\frac{1}{3}\Phi g^{cj}(\nabla_{a}\widehat{\Gamma}_{ci}^{a}-\nabla_{c}\widehat{\Gamma}_{ia}^{a})+\frac{2}{3}\Phi g^{ab}(\nabla_{i}\widehat{\Gamma}^{j}_{ab}-\nabla_{a}\widehat{\Gamma}^{j}_{bi})\\ &+\frac{1}{3}\Phi g^{cj}(\widehat{\nabla}_{a}\widetilde{\Gamma}_{ci}^{a}-\widehat{\nabla}_{c}\widetilde{\Gamma}_{ia}^{a})+\frac{2}{3}\Phi g^{ab}(\widehat{\nabla}_{i}\widetilde{\Gamma}^{j}_{ab}-\widehat{\nabla}_{a}\widetilde{\Gamma}^{j}_{bi})\\ &+\frac{1}{3}\Phi g^{cj}(\widehat{\Gamma}^{a}_{cb}\Gamma^{b}_{ai}-\widehat{\Gamma}^{a}_{ab}\Gamma_{ci}^{b})+\frac{2}{3}\Phi g^{ab}(\widehat{\Gamma}^{c}_{ib}\Gamma_{ac}^{j}-\widehat{\Gamma}^{c}_{ab}\Gamma^{j}_{ci})\\ &+\frac{1}{3}\Phi g^{cj}(\widetilde{\Gamma}^{a}_{cb}\widehat{\Gamma}^{b}_{ai}-\widetilde{\Gamma}^{a}_{ab}\widehat{\Gamma}_{ci}^{b})+\frac{2}{3}\Phi g^{ab}(\widetilde{\Gamma}^{c}_{ib}\widehat{\Gamma}_{ac}^{j}-\widetilde{\Gamma}^{c}_{ab}\widehat{\Gamma}^{j}_{ci})\\ &+\frac{1}{3}\bigl{(}\widehat{\Phi}\widetilde{g}^{cj}+\Phi\widehat{g}^{cj}\bigr{)}(\widetilde{\nabla}_{a}\widetilde{\Gamma}_{ci}^{a}-\widetilde{\nabla}_{c}\widetilde{\Gamma}_{ia}^{a})+\frac{2}{3}\bigl{(}\widehat{\Phi}\widetilde{g}^{ab}+\Phi\widehat{g}^{ab}\bigr{)}(\widetilde{\nabla}_{i}\widetilde{\Gamma}^{j}_{ab}-\widetilde{\nabla}_{a}\widetilde{\Gamma}^{j}_{bi})\\ &+\frac{1}{3}\bigl{(}\widehat{\Phi}\widetilde{g}^{cj}+\Phi\widehat{g}^{cj}\bigr{)}(\widetilde{\Gamma}^{a}_{cb}\widetilde{\Gamma}^{b}_{ai}-\widetilde{\Gamma}^{a}_{ab}\widetilde{\Gamma}_{ci}^{b})+\frac{2}{3}\bigl{(}\widehat{\Phi}\widetilde{g}^{ab}+\Phi\widehat{g}^{ab}\bigr{)}(\widetilde{\Gamma}^{c}_{ib}\widetilde{\Gamma}_{ac}^{j}-\widetilde{\Gamma}^{c}_{ab}\widetilde{\Gamma}^{j}_{ci})\,.\end{split}$ (2.51)
For the equation (2.45), we first consider the contracted second variation equation, obtained by contracting (2.26):
$\partial_{s}k_{l}{}^{l}+\Phi(k_{l}{}^{l})^{2}=\Delta_{g}\phi-\Phi R+3\Lambda\Phi$ (2.52)
and eliminate the scalar curvature using the Hamiltonian constraint (2.14):
@@
$\widetilde{R}+(k_{l}{}^{l})^{2}-\widetilde{k}_{i}{}^{j}\widetilde{k}_{j}{}^{i}=2\widetilde{{\bm{\mathrm{Ric}}}}_{00}+\widetilde{{\bm{R}}}=2\Lambda+\widetilde{\Phi}^{-1}\widetilde{\mathfrak{J}}_{\Phi}-\widetilde{\Phi}^{-1}(\widetilde{\mathfrak{J}}_{k})_{j}{}^{j}$ (2.55)
@@
$\partial_{s}\widehat{k}_{l}{}^{l}=\Delta\Phi-\widetilde{\Delta}\widetilde{\Phi}-\Phi k_{i}{}^{j}k_{j}{}^{i}+\widetilde{\Phi}\widetilde{k}_{i}{}^{j}\widetilde{k}_{j}{}^{i}+\Lambda\widehat{\Phi}+\widetilde{\mathfrak{J}}_{\Phi}$ (2.57)
@@
$\begin{split}\Phi k_{i}{}^{j}k_{j}{}^{i}-\widetilde{\Phi}\widetilde{k}_{i}{}^{j}\widetilde{k}_{j}{}^{i}=&\,\widehat{\Phi}k_{i}{}^{j}k_{j}{}^{i}+\widetilde{\Phi}\widehat{k}_{i}{}^{j}k_{j}{}^{i}+\widetilde{\Phi}\widetilde{k}_{i}{}^{j}\widehat{k}_{j}{}^{i}\\ =&\,\widehat{\Phi}\widehat{k}_{i}{}^{j}\widehat{k}_{j}{}^{i}+2\widehat{\Phi}\widehat{k}_{i}{}^{j}\widetilde{k}_{j}{}^{i}+\widehat{\Phi}\widetilde{k}_{i}{}^{j}\widetilde{k}_{j}{}^{i}+\widetilde{\Phi}\widehat{k}_{i}{}^{j}\widehat{k}_{j}{}^{i}+2\widetilde{\Phi}\widehat{k}_{i}{}^{j}\widetilde{k}_{j}{}^{i}\\ =&\,\Lambda\widehat{\Phi}+2H\widehat{k}_{j}{}^{j}\\ &+\Phi\widehat{k}_{i}{}^{j}\widehat{k}_{j}{}^{i}+2\widehat{\Phi}\widehat{k}_{i}{}^{j}\widetilde{k}_{j}{}^{i}+\widehat{\Phi}\Bigl{(}\widetilde{k}_{i}{}^{j}\widetilde{k}_{j}{}^{i}-\Lambda\Bigr{)}+2\widetilde{\Phi}\widehat{k}_{i}{}^{j}\Bigl{(}\widetilde{k}_{j}{}^{i}-H\widetilde{\Phi}^{-1}\delta_{j}{}^{i}\Bigr{)}\end{split}$ (2.58)
@@
$\begin{split}\Delta\Phi-\widetilde{\Delta}\widetilde{\Phi}=&g^{ij}\nabla_{i}\partial_{j}\Phi+\widehat{g}^{ij}\widetilde{\nabla}_{i}\partial_{j}\widetilde{\Phi}-g^{ij}\widetilde{\nabla}_{i}\partial_{j}\widetilde{\Phi}\\ =&\Delta\widehat{\Phi}-g^{ij}\widehat{\Gamma}_{ij}^{k}\partial_{k}\widetilde{\Phi}+\widehat{g}^{ij}\widetilde{\nabla}_{i}\partial_{j}\widetilde{\Phi}\end{split}$ (2.59)
@@
$\nabla_{j}\widehat{k}_{i}{}^{j}=$ $\,\partial_{i}\widehat{\Phi}-\widehat{\Gamma}_{jc}^{j}(\widetilde{k}_{i}{}^{c}-H\delta_{i}{}^{c})+\widehat{\Gamma}_{ji}^{c}(\widetilde{k}_{c}{}^{j}-H\delta_{c}{}^{j})+\widetilde{\mathfrak{C}}_{i}\,,$ (2.61)
$g^{im}\nabla_{m}\widehat{k}_{i}{}^{j}=$ $\,g^{jc}\partial_{c}\widehat{\Phi}+\widehat{g}^{jc}\widetilde{\nabla}_{c}(\widetilde{k}_{l}{}^{l}-3H)-\widehat{g}^{im}\widetilde{\nabla}_{m}(\widetilde{k}_{i}{}^{j}-H\delta_{i}{}^{j})$ (2.62)
@@
The terms ~$\widetilde{\mathfrak{C}}_{i},\widetilde{\mathfrak{C}}^{j}$ only contain variables of the reference metric $\widetilde{{\bm{g}}}$ and are equal to:
~$\widetilde{\mathfrak{C}}_{i}=-\widetilde{{\bm{\mathrm{Ric}}}}_{0i}\,,\qquad\widetilde{\mathfrak{C}}^{j}=\widetilde{g}^{ij}\widetilde{\mathfrak{C}}_{i}\,.$ (2.63)
@@
~$\widetilde{\nabla}_{j}\widetilde{k}_{i}{}^{j}-\widetilde{\nabla}_{i}\widetilde{k}_{l}{}^{l}={\bm{\mathrm{Ric}}}[\widetilde{{\bm{g}}}]_{0i}$ (2.64)
@@
The equations (2.44), (2.39) are not symmetric hyperbolic in $\widehat{k}_{i}{}^{j},\widehat{\Gamma}_{ic}^{a}$, due to the presence of the terms $\frac{1}{3}\Phi g^{cj}\nabla_{c}\widehat{\Gamma}_{ia}^{a}$, $-\frac{2}{3}\Phi g^{ab}\nabla_{i}\widehat{\Gamma}_{ab}^{j}$ in the RHS of (2.44). However, the latter terms can be treated in the energy estimates by integrating by parts and using the constraint equations (2.61), see Section 5.1.
@@ 3 The background reference metric
$f=\mathcal{O}(\eta e^{mHs})$
for some $m\in\mathbb{Z}$ and $\eta>0$, if $f(s,x)$ is a smooth (analytic) function depending only on the Kerr de Sitter metrics considered, with the property that
@@ 3.1 Kerr de Sitter metric
where we adopt the convention:
$\rho^{2}=r^{2}+a^{2}\cos^{2}\theta,\qquad\Delta_{r}=(r^{2}+a^{2})(1-\frac{\Lambda}{3}r^{2})-2mr,$ (3.3a)
$\Delta_{\theta}=1+\frac{\Lambda}{3}a^{2}\cos^{2}\theta,\qquad\Delta_{0}=1+\frac{\Lambda}{3}a^{2}.$ (3.3b)
The cosmological region is the domain $\Delta_{r}<0$, where $r$ is a time function. With the following reparametrization of the time function, to
$s=H^{-1}\ln r\quad\Leftrightarrow\quad r=e^{Hs}\,,\qquad H=\sqrt{\frac{\Lambda}{3}}\,.$ (3.4)
@@
${\bm{g}}_{\mathcal{K}_{a,m}}=-(\Phi_{\mathcal{K}_{a,m}})^{2}\mathrm{d}s^{2}+(g_{\mathcal{K}_{a,m}})_{ij}\mathrm{d}x^{i}\mathrm{d}x^{j},\qquad x^{1}=t,\,x^{2}=\theta,\,x^{3}=\phi,$ (3.5)
@@
$(g_{\mathcal{K}_{a,m}})_{11}=$ $\,H^{2}e^{2Hs}+\mathcal{O}(1)\,,\qquad(g_{\mathcal{K}_{a,m}})_{13}=\,-\frac{a\sin^{2}\theta}{\Delta_{0}}H^{2}e^{2Hs}+\mathcal{O}(1)\,,$ (3.6)
$(g_{\mathcal{K}_{a,m}})_{22}=$ $\,\frac{e^{2Hs}}{\Delta_{\theta}}+\mathcal{O}(1)\,,\qquad(g_{\mathcal{K}_{a,m}})_{33}=\,\frac{\sin^{2}\theta\Delta_{\theta}+H^{2}a^{2}\sin^{4}\theta}{\Delta_{0}^{2}}e^{2Hs}+\mathcal{O}(1),$
@@
Remark 3.2.
In polar coordinates $(x^{2}=\theta,x^{3}=\phi)$ the metric components degenerate at the poles $\theta=0,\pi$. The expressions (3.1) can then be viewed as the leading order expressions of the metric components in the chart $|\theta-\pi/2|<\pi/3$, $\phi\in(0,2\pi)$. An atlas can be constructed from several of these charts, with the metric components taking identical form in each of them.
@@ 3.2 Partition of Kerr de Sitter
$|a_{1}-a_{2}|+|m_{1}-m_{2}|<\widetilde{\varepsilon}.$ (3.7)
We define a metric $\widetilde{{\bm{g}}}$ as a smooth transition from ${\bm{g}}_{\mathcal{K}_{a_{1},m_{1}}}$ to ${\bm{g}}_{\mathcal{K}_{a_{2},m_{2}}}$:
@@
where $\chi:\mathcal{M}\to[0,1]$ is a smooth function satisfying
$\chi(s,t,\theta,\phi)=\left\{\begin{array}[]{cc}0,&t\leq-1\\ 1,&t\geq 1\end{array}\right.,\qquad|\partial^{\alpha}\chi|\leq C_{\alpha},$ (3.11)
for any coordinate derivative and multi-index $\alpha$. Relative to the coordinates used in the previous subsection (recall Remark 3.2), we have
@@
We call the reference metric (3.8) a partition of Kerr de Sitter. The main properties are recorded in the following Lemma.
@@
$\widetilde{\Phi}-1=$ $\mathcal{O}(e^{-2Hs}),\qquad\widetilde{g}_{ij}=\mathcal{O}(e^{2Hs}),$ (3.13)
$\widetilde{\Gamma}_{ic}^{a}=$ $\mathcal{O}(1),\qquad\widetilde{k}_{i}{}^{j}-H\delta_{i}{}^{j}=\mathcal{O}(e^{-2Hs}),$ (3.14)
@@
$C^{-1}e^{2Hs}\leq\widetilde{g}_{ij}\leq Ce^{2Hs},\qquad|\partial^{\alpha}\widetilde{g}_{ij}|\leq C_{\alpha}e^{2Hs}$
for all $(s,x)\in\mathcal{M}$. Then the statement for $\widetilde{\Gamma}_{ic}^{a}=\frac{1}{2}\widetilde{g}^{al}(\partial_{i}\widetilde{g}_{cl}+\partial_{c}\widetilde{g}_{il}-\partial_{l}\widetilde{g}_{ic})$ becomes obvious.
@@
$2\Phi_{\mathcal{K}_{a,m}}(k_{\mathcal{K}_{a,m}})_{ij}=\partial_{s}(g_{\mathcal{K}_{a,m}})_{ij}=2H(g_{\mathcal{K}_{a,m}})_{ij}+\mathcal{O}(1)$
@@
$2\widetilde{\Phi}\widetilde{k}_{ij}=\partial_{s}\widetilde{g}_{ij}=2H\widetilde{g}_{ij}+\mathcal{O}(1)$

### v2

@@ 2.2.2 First and second variation equations for differences
$\begin{split}\Phi\operatorname{Ric}_{i}{}^{j}-\widetilde{\Phi}\widetilde{\operatorname{Ric}}_{i}{}^{j}=&\phantom{+}\frac{1}{3}\Phi g^{cj}(\nabla_{a}\widehat{\Gamma}_{ci}^{a}-\nabla_{c}\widehat{\Gamma}_{ia}^{a})+\frac{2}{3}\Phi g^{ab}(\nabla_{i}\widehat{\Gamma}^{j}_{ab}-\nabla_{a}\widehat{\Gamma}^{j}_{bi})\\ &+\frac{1}{3}\Phi g^{cj}(\widehat{\nabla}_{a}\widetilde{\Gamma}_{ci}^{a}-\widehat{\nabla}_{c}\widetilde{\Gamma}_{ia}^{a})+\frac{2}{3}\Phi g^{ab}(\widehat{\nabla}_{i}\widetilde{\Gamma}^{j}_{ab}-\widehat{\nabla}_{a}\widetilde{\Gamma}^{j}_{bi})\\ &+\frac{1}{3}\Phi g^{cj}(\widehat{\Gamma}^{a}_{cb}\Gamma^{b}_{ai}-\widehat{\Gamma}^{a}_{ab}\Gamma_{ci}^{b})+\frac{2}{3}\Phi g^{ab}(\widehat{\Gamma}^{c}_{ib}\Gamma_{ac}^{j}-\widehat{\Gamma}^{c}_{ab}\Gamma^{j}_{ci})\\ &+\frac{1}{3}\Phi g^{cj}(\widetilde{\Gamma}^{a}_{cb}\widehat{\Gamma}^{b}_{ai}-\widetilde{\Gamma}^{a}_{ab}\widehat{\Gamma}_{ci}^{b})+\frac{2}{3}\Phi g^{ab}(\widetilde{\Gamma}^{c}_{ib}\widehat{\Gamma}_{ac}^{j}-\widetilde{\Gamma}^{c}_{ab}\widehat{\Gamma}^{j}_{ci})\\ &+\frac{1}{3}\bigl(\widehat{\Phi}\widetilde{g}^{cj}+\Phi\widehat{g}^{cj}\bigr)(\widetilde{\nabla}_{a}\widetilde{\Gamma}_{ci}^{a}-\widetilde{\nabla}_{c}\widetilde{\Gamma}_{ia}^{a})+\frac{2}{3}\bigl(\widehat{\Phi}\widetilde{g}^{ab}+\Phi\widehat{g}^{ab}\bigr)(\widetilde{\nabla}_{i}\widetilde{\Gamma}^{j}_{ab}-\widetilde{\nabla}_{a}\widetilde{\Gamma}^{j}_{bi})\\ &+\frac{1}{3}\bigl(\widehat{\Phi}\widetilde{g}^{cj}+\Phi\widehat{g}^{cj}\bigr)(\widetilde{\Gamma}^{a}_{cb}\widetilde{\Gamma}^{b}_{ai}-\widetilde{\Gamma}^{a}_{ab}\widetilde{\Gamma}_{ci}^{b})+\frac{2}{3}\bigl(\widehat{\Phi}\widetilde{g}^{ab}+\Phi\widehat{g}^{ab}\bigr)(\widetilde{\Gamma}^{c}_{ib}\widetilde{\Gamma}_{ac}^{j}-\widetilde{\Gamma}^{c}_{ab}\widetilde{\Gamma}^{j}_{ci})\end{split}$ (2.48)
For the equation (2.42), we consider (2.16) for a solution to (1.1):
@@
$\widetilde{R}+(\widetilde{k}_{l}{}^{l})^{2}-\widetilde{k}_{i}{}^{j}\widetilde{k}_{j}{}^{i}=2\widetilde{{\bm{\mathrm{Ric}}}}_{00}+\widetilde{{\bm{R}}}=2\Lambda+\widetilde{\Phi}^{-1}\widetilde{\mathfrak{J}}_{\Phi}-\widetilde{\Phi}^{-1}(\widetilde{\mathfrak{J}}_{k})_{j}{}^{j}$ (2.51)
@@
$\partial_{s}\widehat{k}_{l}{}^{l}=\Delta\Phi-\widetilde{\Delta}\widetilde{\Phi}-\Phi k_{i}{}^{j}k_{j}{}^{i}+\widetilde{\Phi}\widetilde{k}_{i}{}^{j}\widetilde{k}_{j}{}^{i}+\Lambda\widehat{\Phi}+\widetilde{\mathfrak{J}}_{\Phi}\,.$ (2.53)
@@
[…]
@@
[…]
@@
$\nabla_{j}\widehat{k}_{i}{}^{j}$= $\,\partial_{i}\widehat{\Phi}-\widehat{\Gamma}_{jc}^{j}(\widetilde{k}_{i}{}^{c}-H\delta_{i}{}^{c})+\widehat{\Gamma}_{ji}^{c}(\widetilde{k}_{c}{}^{j}-H\delta_{c}{}^{j})+\widetilde{\mathfrak{C}}_{i}$ (2.57)
[…]
@@
The terms $\widetilde{\mathfrak{C}}_{i},\widetilde{\mathfrak{C}}^{j}$ only contain variables of the reference metric $\widetilde{{\bm{g}}}$ and are equal to:
$\widetilde{\mathfrak{C}}_{i}=-\widetilde{{\bm{\mathrm{Ric}}}}_{0i}\,,\qquad\widetilde{\mathfrak{C}}^{j}=\widetilde{g}^{ij}\widetilde{\mathfrak{C}}_{i}\,.$ (2.59)
@@
$\widetilde{\nabla}_{j}\widetilde{k}_{i}{}^{j}-\widetilde{\nabla}_{i}\widetilde{k}_{l}{}^{l}={\bm{\mathrm{Ric}}}[\widetilde{{\bm{g}}}]_{0i}\,$, (2.60)
@@
The equations (2.41), (2.36) are not symmetric hyperbolic in $\widehat{k}_{i}{}^{j},\widehat{\Gamma}_{ic}^{a}$, due to the presence of the terms $\frac{1}{3}\Phi g^{cj}\nabla_{c}\widehat{\Gamma}_{ia}^{a}$, $\frac{2}{3}\Phi g^{ab}\nabla_{i}\widehat{\Gamma}_{ab}^{j}$ in the RHS of (2.41). However, the latter terms can be treated in the energy estimates by integrating by parts and using the constraint equations (2.57), see Section 5.1.
The system of equations derived in this section are summarized in Appendix B.
@@ 3 The background reference metric
$f=\mathcal{O}(\eta e^{mHs})\,$,
for some $m\in\mathbb{Z}$ and $\eta>0$, if $f(s,x)$ is a smooth function depending only on the Kerr de Sitter metrics considered, with the property that
@@ 3.1 Kerr de Sitter metric
where we adopt the convention [Car09]:
$\rho^{2}=r^{2}+a^{2}\cos^{2}\theta,\hskip 18.49988pt\Delta_{r}=(r^{2}+a^{2})(1-\frac{\Lambda}{3}r^{2})-2mr,$ (3.3a)
$\Delta_{\theta}=1+\frac{\Lambda}{3}a^{2}\cos^{2}\theta,\hskip 18.49988pt\Delta_{0}=1+\frac{\Lambda}{3}a^{2}.$ (3.3b)
The cosmological region is the domain $\Delta_{r}<0$, where $r$ is a time function. With the following reparametrization of the time function $r$, to
$s=H^{-1}\ln r\qquad\Leftrightarrow\qquad r=e^{Hs}\,,\hskip 18.49988ptH=\sqrt{\frac{\Lambda}{3}}\,$, (3.4)
@@
${\bm{g}}_{\mathcal{K}_{a,m}}=-(\Phi_{\mathcal{K}_{a,m}})^{2}\mathrm{d}s^{2}+(g_{\mathcal{K}_{a,m}})_{ij}\mathrm{d}x^{i}\mathrm{d}x^{j},\hskip 18.49988ptx^{1}=t,\,x^{2}=\theta,\,x^{3}=\phi\,$, (3.5)
@@
$(g_{\mathcal{K}_{a,m}})_{11}=$ $\,H^{2}e^{2Hs}+\mathcal{O}(1)\,,\hskip 18.49988pt(g_{\mathcal{K}_{a,m}})_{13}=\,-\frac{a\sin^{2}\theta}{\Delta_{0}}H^{2}e^{2Hs}+\mathcal{O}(1)\,$, (3.6)
$(g_{\mathcal{K}_{a,m}})_{22}=$ $\,\frac{e^{2Hs}}{\Delta_{\theta}}+\mathcal{O}(1)\,,\hskip 18.49988pt(g_{\mathcal{K}_{a,m}})_{33}=\,\frac{\sin^{2}\theta\Delta_{\theta}+H^{2}a^{2}\sin^{4}\theta}{\Delta_{0}^{2}}e^{2Hs}+\mathcal{O}(1)\,.$
@@
This transformation is motivated by the standard form of FLRW spacetimes; see [Fou22].
@@ 3.2 Partition of Kerr de Sitter
The Kerr de Sitter metrics ${\bm{g}}_{\mathcal{K}_{a,m}}$ can be viewed as a 2-parameter family of metrics on the fixed differentiable structure of the cosmological region of a Schwarzschild de Sitter spacetime ${\bm{g}}_{m,0}$. In other words, we can view $({\bm{g}}_{\mathcal{K}_{a,m}})_{ss}$, and $({\bm{g}}_{\mathcal{K}_{a,m}})_{ij}$ as a family of metric components in a fixed chart
$(s,t,\phi,\theta)\in(s_{\mathcal{C}},\infty)\times\mathbb{R}\times(0,2\pi)\times(0,\pi)\,.$ (3.7)
@@
$|a_{1}-a_{2}|+|m_{1}-m_{2}|<\widetilde{\varepsilon}\,$, (3.8)
and choose $m$ any value between $m_{1}$ and $m_{2}$.
We define a metric $\widetilde{{\bm{g}}}$ in the above chart as a smooth transition from ${\bm{g}}_{\mathcal{K}_{a_{1},m_{1}}}$ to ${\bm{g}}_{\mathcal{K}_{a_{2},m_{2}}}$:
@@
where $\chi:\mathcal{R}\to[0,1]$ is a smooth function represented by
$\chi(s,t,\theta,\phi)=\left\{\begin{array}[]{cc}0,&t\leq-1\\ 1,&t\geq 1\end{array}\right.,\hskip 18.49988pt|\partial_{\alpha}\chi|\leq C_{\alpha},$ (3.12)
for any coordinate derivative and multi-index $\alpha$. We call the reference metric (3.9) a partition of Kerr de Sitter. In the fixed coordinate chart we have
@@
Remark 3.2.
In view of the use of polar coordinates on the cylinder $(x^{1},x^{2},x^{3})=(t,\theta,\phi)$ the spacetime manifold is not covered by a single chart (3.7). However, an atlas can be constructed from two of these charts. For the pointwise bounds below it is understood that the chart is restricted away from the points where the metric degenerates in these coordinates.
The main properties are recorded in the following Lemma.
@@
$\widetilde{\Phi}-1=\mathcal{O}(e^{-2Hs}),\hskip 18.49988pt\widetilde{g}_{ij}=\mathcal{O}(e^{2Hs}),\hskip 18.49988pt\widetilde{g}^{ij}=\mathcal{O}(e^{-2Hs}),$ (3.14)
$\widetilde{\Gamma}_{ic}^{a}=\mathcal{O}(1),\hskip 18.49988pt\widetilde{k}_{i}{}^{j}-H\delta_{i}{}^{j}=\mathcal{O}(e^{-2Hs}),$ (3.15)
@@
$C^{-1}e^{2Hs}\leq\widetilde{g}_{ij}\leq Ce^{2Hs},\hskip 18.49988pt|\partial^{\alpha}\widetilde{g}_{ij}|\leq C_{\alpha}e^{2Hs}\,$,
for all $(s,x)\in\mathcal{M}$. This implies both the statements for $\widetilde{g}_{ij}$ and $\widetilde{g}^{ij}$. Then the statement for $\widetilde{\Gamma}_{ic}^{a}=\frac{1}{2}\widetilde{g}^{al}(\partial_{i}\widetilde{g}_{cl}+\partial_{c}\widetilde{g}_{il}-\partial_{l}\widetilde{g}_{ic})$ becomes obvious.
@@
$2\Phi_{\mathcal{K}_{a,m}}(k_{\mathcal{K}_{a,m}})_{ij}=\partial_{s}(g_{\mathcal{K}_{a,m}})_{ij}=2H(g_{\mathcal{K}_{a,m}})_{ij}+\mathcal{O}(1)\,$,
@@
$2\widetilde{\Phi}\widetilde{k}_{ij}=\partial_{s}\widetilde{g}_{ij}=2H\widetilde{g}_{ij}+\mathcal{O}(1)\,$,